what is the standard form of y^2-2y-4x-31=0
Answer by lwsshak3(11628)

what is the standard form of y^2-2y-4x-31=0
**
y^2-2y-4x-31=0
complete the square
(y^2-2y+1)-4x-31-1=0
(y-1)^2-4x-32=0
(y-1)^2-4(x+8)=0
(y-1)^2=4(x+8)
This is a standard form of equation for a parabola with vertex at (-8,1)
Parabola opens rightwards with axis of symmetry at y=1